# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ 5.3 Possible explanations of the long-term radio flare
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997). Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that the advection-dominated accretion flows cannot explain the observed radio emission in our case.
@@
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG; Green et al. 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993). Later, Behar et al. (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. Magnetized coronal winds were proposed to explain the observed radio emission of the Seyfert galaxy Mrk 590 by Koay et al. (2016), however, the authors noted that a jet contribution cannot be ruled out. These two types of outflows might even be equivalent and cannot be distinguished in Seyfert galaxies with unresolved compact core features seen in their VLBI maps (e.g., Kharb et al. 2015), like in the case of AT2019aalc. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be rapidly variable at radio wavelengths and should evolve similarly to the X-ray flux, which is clearly not the case. Although, a contribution from the mechanisms discussed so far to the total radio flux remains a possibility, we need to look for another possible origin to explain the main origin.
@@
Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets which can be either on or off-axis relative to the observer. These features are often visible on VLBI radio maps of radio-detected Seyfert galaxies (e.g., Kozák et al. 2024) and a few TDEs (e.g., Paragi et al. 2017; Mohan et al. 2022). In our case, the high brightness temperature of the EVN-detected radio feature together with its position (in good agreement with the optical Gaia position) indicate its radio core identification, however, no other components were detected. Nevertheless, the steep () radio spectrum revealed by our ATCA observations indicates the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum. The steep radio spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s). The ATCA radio spectrum indicates a high-frequency turnover above GHz which is a possible hint of a newly ejected radio-emitting component. We further discuss the radio spectrum of AT2019aalc in the following.
@@
The turnover suggests the multiple component nature of the spectrum. Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al. 2020). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. Since older sources are larger than young ones, the expanding relativistic jet or non-relativistic outflow loses energy as the density decreases and the spectral peak moves to lower energies over time. With other words, as the projected linear size of the radio source increases, the turnover frequency decreases. These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al. 2020). The low-frequency peak is, however, more likely to be explained with star formation activity (as NLSy1 galaxies have a significantly higher star formation rate as other AGNs) instead of AGN emission. The latter is assumed to be synchrotron self-absorbed (SSA) to explain the faintness or the non-detection of these galaxies at lower frequencies. As another possibility to explain the high-frequency excess of, Berton et al. (2020) discusses a scenario in which the AGN emission is absorbed below GHz due to free-free absorption (FFA). In that case, the circumnuclear gas might be ionized both by the AGN and by hot stars and might act as a screen for the jet at low frequencies or the passage of the jet in the interstellar medium causes the formation of a cocoon of ionized gas and this way being responsible for the FFA. Finally, coronal emission could also be responsible for the inverted spectrum.
@@
In the case of AT2019aalc, the coronal emission scenario can be safely excluded as we should see a rapidly variable radio light curve which should evolve similarly to the X-ray emission. On the other hand, the other two scenarios might also be challenging. The large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA. Moreover, the presence of a relatively strong [O iii] line in the SDSS host spectrum implies earlier AGN activity as being responsible for the radio emission before the long-term flare. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
Nevertheless, the contribution of a newly ejected component to the overall radio spectrum remains a possibility. In the cases of the earlier discussed NLSy1 galaxies, the high frequency peak is thought to be suggesting a newly born radio source with no signs of earlier jet activity. Similarly to gigahertz-peaked sources (GPS), the high-frequency peak is thought to be shifting towards lower energies over time. This way, NLSy1 galaxies might be extremely young GPS which was first suggested by Oshlack et al. (2001). In our case, a newly ejected component can be responsible for the high-frequency excess whilst the low-frequency part is naturally explained with optically thin synchrotron emission of the AGN indicating that we do not see a newly born radio-AGN but an AGN with recently enhanced, intermittent activity. The result is a two-component radio spectrum with a relatively steep spectrum at lower frequencies and a high-frequency excess. The steepness of the spectrum between MHz and GHz might be explained with cooling of electrons originating from past AGN activity, and also a contribution from star formation can play a role here. Järvelä et al. (2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al. (2020). The broad-line AGN classification of the host galaxy of AT2019aalc is more consistent with a restarted AGN picture as this type of AGN are thought to be older than NLSy1s making a restarted activity scenario over a newly born radio source one more probable. The restarted activity results in a peaked spectrum at higher frequencies, superimposed on a spectrum from a period (or periods) of earlier activity peaking at lower frequencies. Intermittent activity is further supported by theoretical models suggesting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al. 2009). This is compatible with the accretion-induced lines seen in the post-flare-1 optical spectrum of AT2019aalc (see Subsubsect. 4.5.2 and van Velzen et al. 2024).
We note that the fitted ellipsoid model fit component to the EVN visibility data is aligned to the northwest-southeast direction (clearly differs from the position angle of the restoring beam indicating the angle-dependent resolution of the VLBI array) which potentially indicates a newly ejected radio-emitting component. Higher frequency observations with improved angular resolution are required in order to resolve the radio structure we detected at GHz. We observed AT2019aalc with the EVN and the Korean VLBI Network (KVN) jointly at GHz in March and expect to reach sub-mas-scale angular resolution, roughly an order of magnitude better than with the currently discussed observation. These results will be presented in a follow-up publication
@@ 5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al. 2015). Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess. This soft excess can be well modeled by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al. 2012).
The stacked Swift/XRT X-ray spectrum of AT2019aalc can be adequately described with an absorbed blackbody+power-law model. We note that in the cases of strongly beamed jetted sources e.g. blazars, the X-ray spectrum can be well modeled with a simple or broken power-law without adding any additional model component (e.g., Kammoun et al. 2015; Acciari et al. 2022). Although AT2019aalc shows enhanced radio activity, its brightness temperature suggests a non-beamed nature of its radio emission in agreement with the X-ray spectrum.
@@
The X-ray spectrum is dominated by a power-law component with a soft power-law index of while the thermal component is extremely soft with a blackbody temperature of eV derived from the Swift/XRT observations. Most of the BLSy1s typically have a significantly harder X-ray spectrum with a power-law index (e.g., Pons & Watson 2014). Our value is more typical for sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al. 1996; Saxton et al. 2020).
The soft X-ray excess appears in the stacked spectrum and becomes more significant over time as we added a blackbody component in order to better describe the X-ray spectrum. Furthermore, the data of the peak of the second X-ray flare is clearly better-fit with a single blackbody model than with a power-law one (Subsect. 4.3) implying not just the need of the blackbody component to adequately fit the data but even its dominance over the power-law component at this epoch. A softening of the power-law index over time can be seen as well. These results suggest significantly enhanced accretion or disk instability. Delayed accretion flares have been observed in TDEs (Saxton et al. 2020). In our case, the post-flare-1 spectrum shows clear signs of accretion and the second flare is followed by enhanced accretion activity as well.
@@
The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x]. The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al. 2009). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability we consider only observations taken on the same day). This implies that the expected soft X-ray luminosity is consistent with the observed values. The soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as it was suggested for ECLEs by Wang et al. (2012). The photoionization or collisional excitation of the dust grains is rather due to a current event. This further supports that the reoccurring optical/UV outbursts seen during the decaying of the flares are driven by the soft X-ray flares which lead these outbursts only with a couple of weeks. These suggest an accretion-driven origin of the reoccurring outbursts in the optical/UV during the decaying of the flare.
@@ 5.5 SED fitting
The continuum emission of TDEs are well described by a thermal blackbody model (Gezari 2021). The case of AT2019aalc appears to be more complex. An excess in the UW2 band can be seen regardless the consideration of a simple or the double blackbody fit. The UV emission have a contribution from the Bowen Fluorescence mechanism and potentially also from Comptonization explaining the poor fit towards the blue. The red wing of the optical and the IR part cannot be well fitted with a simple blackbody model suggesting a complex, multi-component dusty reprocessing region. The excess in the -band might be explained with the early appearance of some Fe lines in this range (such as [Fe x] and [Fe vii]) indicating the release of this element previously locked up by the dust grains. Since all of this seem to be started already around the optical peak, at least a fraction of the dust is located on subpc scales to the SMBH. This is further consistent with the extreme line widths of the coronal lines (see Subsect. 5.1.3). Therefore, not only the dusty tori is responsible for the IR dust echo. It is possible that the significant amount of dust in the nucleus region is the consequence of enhanced star formation sometime in the past. The Baldwin, Phillips& Terlevich (BPT, Baldwin et al. 1981) emission-line ratio diagnostics of the host SDSS spectrum are log([O iii] and log([N ii] . We plotted the BPT line-ratio of SDSS spectra with indicating the ratios of the host galaxy of AT2019aalc and the BFFs AT2021loi, AT2017bgt and F01004-2237 in Fig. 22 using astroml (Vanderplas et al. 2012; Ivezić et al. 2014). All of the investigated galaxies are in the composite region, consequently, in addition to AGN activity, star formation activity might also have a remarkable contribution for producing the narrow lines seen in the spectra. Moreover, these ratios place these two galaxies solidly in the region of the diagram where the post-starburst sample of French et al. (2015) is located. We note that to derive the ratios for AT2019aalc and F01004-2237, we used pre-flare spectra. While as no host spectra are available for AT2017bgt and AT2021loi, we considered the spectra taken as early as possible in these cases.
@@ 5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021). Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the decay of the optical emission, which is significantly slower than for any TDE identified so far. The color evolution of the flares is also atypical for TDEs.
@@
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of strong high-ionization coronal lines.
The mechanism that powers the long-term accretion enhancement in BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al. 2017) or the AGN’s radiation field. The bumps in optical and UV during the decay of the flares of AT2019aalc led by soft X-ray flares indicate the flare’s accretion-driven origin consistently within this picture. Moreover, the flaring episodes in X-rays seem to be happen periodically which is expected from interaction between a circularizing stellar stream and the accretion disk. The origin through reprocessing is further consistent with the ionization of the BLR and so on the increase in the Bowen and the Balmer lines seen in the spectra during the second flare. The dusty region (expected from a source with IR dust echo) and the liberation of Fe and S from these dust grains due to photoionization by the soft X-ray flares provide an explanation for the high-ionization coronal lines. The increase of the W1W2 color implies significant cooling expected from cooling dust after the onset of the dust echo. This scenario provides explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. The second flare is even more luminous than the first one, similarly to the case of the partial TDE AT2020vdq (Somalwar et al. 2023). A partial TDE scenario might explain the second optical flare.
@@
Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well. The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understood their unique optical light curves. Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs. This would explain the TDE-like properties of these sources. The most remarkable differences i.e. the slowly decaying optical emission and the continuum emission not being well described by blackbody models might appear due to a more complex environment the TDE takes place with comparing to quiescent galaxies. Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al. 2012) and the BFFs is consistent with this picture. A known potential partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al. 2015). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 17). Interestingly, IC 3599 is the only known AGN with fading coronal lines (Frederick et al. 2019).
Alternatively, a double TDE i.e. the tidal disruptions of a pair of binary stars might explain the double-peaked optical light curve. Notably, double TDEs may represent nearly of all stellar tidal disruptions (Mandel & Levin 2015).
The strength of the high-ionozation coronal lines that regular AGN activity cannot explain the unusual behaviour of this transient. It is not clear why we do not see as strong high-ionozation coronal lines in normal AGNs as in the transients with significant dust echo emission i.e. TDEs and BFFs. It might be explained with dust sublimation due to the UV radiation of the AGN (the sublimation radius is typically in the subpc range, van Velzen et al. 2021b) which clears the dust from the nucleus region. AGN-driven outflows play also a role here. The limited variability of the host galaxy of AT2019aalc prior the first flare and its location in the BPT diagram suggest weak AGN activity before the transient occurred. Therefore, the dust possibly originating from star formation activity in the nucleus could survive. Additionally, the first flare could not clear the nucleus region which suggests that it must have been way less luminous in UV than the second flare. Unfortunately, due to lack of UV data around the first flare this hypothesis cannot be further tested. Nevertheless, the detection of some Fe lines in the post-flare-1 spectrum implies that a certain fraction of the dust was destroyed following the first flare.
Interestingly, the host AGN of other BFFs also showed weaker pre-transient activity than persistently accreting AGN (Trakhtenbrot et al. 2019; Makrygianni et al. 2023). The TDE-like features suggest that TDEs might restarted these AGN. Beyond the TDE-related scenarios, it is possible that the restarted AGN activity is rather explained with the dusty environment itself.
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs available. Based on the radio observations explained in Makrygianni et al. (2023) AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz which can be compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The host galaxy of AT2017bgt was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. In addition, the BFF candidate VT J1548 was also found to exhibit a radio spectrum peaking around GHz (in contrast with non-jetted AGN that peak below GHz, e.g., Padovani et al. 2017) which was explained with an ejection of a synchrotron-emitting outflow (Somalwar et al. 2022). These findings allow us to conclude that newly ejected radio outflow/jet components play a crucial role in BFFs. These outflows ejected due to enhanced accretion might be partly responsible for the liberation of Fe and S from the dust grains located relatively close to the SMBH. The collisional ionization in the case of AT2019aalc is supported by the extreme broadness of the coronal lines especially [S xii.

### v2

@@ 5.3.1 Possible explanations of the long-term radio flare
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997). Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
@@
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG: Green et al. 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993). Later, Behar et al. (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be highly variable at radio wavelengths.
@@
Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., Paragi et al. 2017; Mohan et al. 2022; Kozák et al. 2024). The high brightness temperature of the EVN-detected radio feature indicates its radio core identification. The steep () radio spectrum indicate the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum. The spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s). The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
5.3.2 The radio spectrum
@@
The turnover suggests the multiple component nature of the spectrum. Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al. 2020). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al. 2020). The low-frequency peak is, however, more likely to be explained with star formation activity instead of AGN emission since the latter is absorbed due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
@@
However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
A newly ejected component can explain the high-frequency excess whilst the low-frequency part is originating from optically thin synchrotron emission of the AGN. Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior. Järvelä et al. (2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al. (2020). Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al. 2009).
@@ 5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al. 2015). Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess. It can be well modeled by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al. 2012).
@@
The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV. Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., Pons & Watson 2014). Our value is more commonly seen in sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al. 1996; Saxton et al. 2020).
@@
The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x]. The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al. 2009). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day). Thus, the soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as suggested for ECLEs by Wang et al. (2012).
The reoccurring optical/UV bumps seen during the decaying of the main optical flares are likely driven by the soft X-ray flares which lead these outbursts only with a couple of weeks. The soft X-ray excess appears to be more dominant during these flares. Interestingly, the soft X-ray flares show hints of quasi periodicity of days. A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit with soft X-ray flares on time-scales of hours to days while partial TDEs on timescales of months to years (e.g., Chakraborty et al. 2021; Quintin et al. 2023; Miniutti et al. 2023; Nicholl et al. 2024). We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring. This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: Evans et al. 2023) as LSXPS J031335.6020905. The potential connection between BFFs and QPEs should be explored in future studies.
@@ 5.5 SED fitting
The continuum emission of TDEs is well described by a thermal blackbody model (Gezari 2021). The case of AT2019aalc appears to be more complex. The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization explaining the poor fit towards the blue. The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations as the IR light curve is peaked well after the optical one. The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range but such a significant excess cannot be fully explained with the coronal lines only.
The inferred blackbody temperature and radius values estimated for the ’blue’ blackbody fit are K and cm, respectively. The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by van Velzen et al. (2020) and Thomsen et al. (2022). On the other hand, the radius is smaller than calculated for optical selected TDEs but is more consistent with the X-ray detected ones. The ’red’ blackbody fit resulted in K and cm.
@@ 5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021). Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far. The reddening in color when the main optical flares are declining is also atypical for TDEs.
@@
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
The mechanism that powers the BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al. 2017, and ASASSN-14ko, Huang et al. 2023) or the AGN’s radiation field. According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (Chan et al. 2019). The debris-disk interaction results in enhanced accretion and multiple shocks (Chan et al. 2019), notably leading to X-ray emission (Ryu et al. 2024). Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk. These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare. The soft X-ray emission is likely reprocessed in optical/UV by the dense environment explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines. The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (Chan et al. 2019; Ryu et al. 2024). Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (Somalwar et al. 2023) and the BFF F01004-2237. Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (Zhong 2024; Liu et al. 2024). Liu et al. (2024) claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq. A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2017bgt and AT2021loi) with double-peaked optical flares as well.
@@
Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well. The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understand their unique optical light curves. Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs. This would explain the TDE-like properties of these sources. The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”). TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (Chan et al. 2019). Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al. 2012) and the BFFs (see Fig.13) is consistent with this picture. A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al. 2015). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 13). In addition, IC 3599 is the only known AGN with fading coronal lines (Frederick et al. 2019).
5.7 Radio properties of BFFs
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs. Based on the radio observations explained in Makrygianni et al. (2023), AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. The BFF AT2020afhd was detected with the VLA at GHz with mJy (Christy et al. 2024). If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations published in Chan et al. (2019). Radio observations of any newly discovered BFF will be essential to better understand these transients.